# IEPile: Unearthing Large-Scale Schema-Based Information Extraction Corpus

### v2

@@ Abstract
Large Language Models (LLMs) demonstrate remarkable potential across various domains; however, they exhibit a significant performance gap in Information Extraction (IE). Note that high-quality instruction data is the vital key for enhancing the specific capabilities of LLMs, while current IE datasets tend to be small in scale, fragmented, and lack standardized schema. To this end, we introduce IEPile, a comprehensive bilingual (English and Chinese) IE instruction corpus, which contains approximately 0.32B tokens. We construct IEPile by collecting and cleaning 33 existing IE datasets, and introduce schema-based instruction generation to unearth a large-scale corpus. Experimental results on LLaMA, Baichuan and Qwen demonstrate that using IEPile can enhance the performance of LLMs for IE, especially the zero-shot generalization. We open-source the resource and pre-trained models, hoping to provide valuable support to the NLP community.
@@
Honghao ${}^{\spadesuit}$${}^{\diamondsuit}$thanks:     Equal Contribution., Lin ${}^{\clubsuit}$${}^{\diamondsuit}$11footnotemark: 1, Hongbin ${}^{\spadesuit}$, Ningyu ${}^{\spadesuit}$${}^{\diamondsuit}$thanks:     Corresponding Author. Mengshu ${}^{\clubsuit}$${}^{\diamondsuit}$, Lei ${}^{\clubsuit}$${}^{\diamondsuit}$, Huajun ${}^{\spadesuit}$${}^{\diamondsuit}$22footnotemark: 2 ${}^{\spadesuit}$ Zhejiang University ${}^{\diamondsuit}$ Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph ${}^{\clubsuit}$ Ant Group {guihonghao,zhangningyu,huajunsir}@zju.edu.cn https://github.com/zjunlp/IEPile
@@ 1 Introduction
By fine-tuning the Baichuan (Yang et al., 2023), LLaMA (Touvron et al., 2023b), and Qwen (Bai et al., 2023), we show that LLMs with IEPile can yield better zero-shot performance than baselines. This achievement not only verifies the effectiveness of the IEPile dataset but also provides a framework for creating IE datasets in other domains.
@@ 2.1 Data Collection and Cleaning
To broadly cover various domains and meet the practical demands, we collect datasets necessary for IE from multiple data sources. Our corpus mainly involves bilingual data (Chinese and English) and focuses on three main categories of IE tasks: Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE). In total, we gather 26 English datasets and 7 Chinese datasets. We also employ standardization procedures to maintain data quality and format uniformity, involving format unification, instance deduplication, and the exclusion of low-quality data.
@@ 2.2 Schema-Based Instruction Generation
We concentrate on instruction-based information extraction (IE), which relies on three elements to compose an instruction: 1) Task Description, a template utilized to distinguish between different IE tasks; 2) Input Text, the source text that the model is expected to extract information; and 3) Schema sequence, which defines the information that the model is supposed to extract, including entity types, relations, events, etc. Among these, the schema sequence is critical as it reflects the specific extraction requirements and is dynamically variable. Therefore, constructing the schema sequence in an instruction is of utmost importance.
@@
Firstly, we define schemas that actually exist within the input text as positive schemas and those that do not appear as negative schemas. As illustrated in Figure 1, the “location contains” present in the annotation is a positive schema, while all other schemas from the predefined label set $L$ are negative schemas. Traditional IE frameworks, which are treated as sequence labeling tasks, take text as input and produce a label for each token as output, without involving the concept of positive or negative schemas within the model’s input. However, in the era of generative IE, represented by models like UIE  (Lu et al., 2022a), introduce the concept of integrating a schema sequence (refers to as Structural Schema Instructor, or SSI) in the model’s input to guide its output, restricting the range of output to the SSI. The concept necessitates including the entire predefined label set of a dataset as the SSI to guide the model’s output during inference. Consequently, if the SSI during the training contains only positive schemas, the model will tend to generate corresponding answers for every label within the SSI during inference. Therefore, to make the model explicitly reject generating outputs for negative schemas, it is necessary to incorporate negative schemas into the SSI.
@@
As illustrated in Figure 1, assume that dataset $\mathcal{D}$ possesses a predefined label set $L$. For a given text $S$, the schemas present in its annotation constitute the positive schema set $Pos\_L$, while others form the negative schema set $Neg\_L$. In our analysis, we discover that the primary cause of model mistakes stems from the semantic ambiguity of the schema. In traditional approaches, the $Neg\_L$ is simply defined as $L-Pos\_L$. However, they overlook a critical aspect: it is important to pay special attention to negative schemas that are semantically close to positive schemas. Inspired by the theory of contrastive learning, we construct a hard negative schema dictionary $\mathcal{K}$, where each key represents a unique schema and the associated value is a collection of schemas that are semantically similar to the key schema. Based on this, we define the hard negative schema set as $Hard\_L=\mathcal{K}[Pos\_L]$, and the other negative schema set as $Other\_L=L-Pos\_L-Hard\_L$. The final $Neg\_L$ is constituted by $Hard\_L$ and a small subset of $Other\_L$. Through this strategy, we not only present semantically similar schemas more frequently within the instruction but also reduce the number of training instances without sacrificing model performance.
@@
Figure 2: Distribution of different tasks, domains, and datasets within IEPile.
@@ 2.3 Data Statistics
Based on the aforementioned methods, we obtain IEPile, which includes roughly 2 million instruction entries and approximately 0.32B tokens (utilizing the Baichuan2 tokenizer). Figure 2 displays the distribution of domains and datasets within IEPile, including 33 datasets spanning multiple domains such as general, news, finance, and biomedical. Additionally, Table B.2 provides examples of instructions and outputs for 3 different tasks within the IEPile dataset.
@@
3 Experiments
Based on IEPile, we fine-tune (Lora) models such as Baichuan2, LLaMA2, and Qwen1.5, subsequently comparing their zero-shot generalization capabilities against a range of baseline models. Results of the full supervision evaluation, and training details are in Appendix C.
3.1 Experimental Settings
Evaluation Metrics: We employ span-based Micro-F1 as the metric for measuring model performance. Baselines: We select a range of strong models for comparative analysis, which include UIE  (Lu et al., 2022a), LLaMA2-13B-Chat, Baichuan2-13B-Chat, Mistral-7B-Instruct-v0.2  (Jiang et al., 2023), Qwen1.5-14B-Chat, ChatGPT  (Ouyang et al., 2022), GPT-4  (OpenAI, 2023), InstructUIE  (Wang et al., 2023b), and YAYI-UIE  (Xiao et al., 2023). Zero-shot Benchmark: We collect 13 datasets that are not present in the training set.
3.2 Main Results
In Tables 1 and  2, we report the zero-shot performance across three tasks: NER, RE, and EE for two languages. Overall, after training with the IEPile, the models achieve better results in the majority of tasks. We believe the success is due to the hard negative schema construction and batched instruction generation strategy, which can mitigate the train-eval mismatch and semantic ambiguity for the diverse schema. We also observe that Baichuan2-IEPile, LLaMA2-IEPile and Qwen1.5-IEPile are slightly behind GPT-4 in English NER. We hypothesize that the marginal gap may be attributed to GPT-4’s exposure to a vast corpus of similar data during its training. Moreover, it is essential to note that InstructUIE focuses on English data while IEPile incorporates both English and Chinese data. This disparity in data may influence the capability of the model in English, potentially reducing the performance.
Inconsistency in the Number of Schema Queries Hurt Generalization.
3.3 Analysis
@@ 4 Conclusion and Future Work
From the data perspective, our study primarily focuses on schema-based IE, which limits our ability to generalize to human instructions that do not follow our specific format requirements. Additionally, we do not explore the field of Open Information Extraction (Open IE); however, if we remove schema constraints, our dataset would be suitable for Open IE scenarios. Besides, IEPile is confined to data in English and Chinese, and in the future, we hope to include data in more languages. From the model’s perspective, our research evaluates only two models, Baichuan and LLaMA, along with a few baselines due to the computation resources. IEPile can be applied to any other LLMs such as ChatGLM Du et al. (2022) and Gemma [^3] https://huggingface.co/google/gemma-7b..
@@ A.1 Information Extraction Datasets
Large-scale pre-trained corpora are crucial for the effectiveness of LLMs, providing a wealth of knowledge and a foundation for language comprehension, while the annotated data for information extraction (IE) also holds its importance. Although the field of IE has accumulated a considerable amount of annotated data (Walker et al., 2006; Riedel et al., 2010; Sang and Meulder, 2003; Luan et al., 2018), these datasets are often limited in size, scattered in distribution, and lack standardization in schema. Faced with these limitations, there is an urgent need for generating instruction data through unified and automated methods to fill the gap of the current lack of centralized, large-scale IE instruction tuning datasets. We focus on instruction-based IE scenarios and, by collecting and cleaning existing IE data, build a large, schema-rich IE instruction dataset called IEPile. IEPile is designed to enhance the adaptability and processing capabilities of LLMs for different IE tasks, simultaneously strengthening their generalization skills to extract from new domains and schemas.
@@ A.2 Information Extraction Models
Recently, LLMs (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023a, b) demonstrate their exceptional versatility and generalization capabilities across a variety of downstream tasks  (Vilar et al., 2023; Hegselmann et al., 2023). Particularly in the domain of IE, these models have the potential to tackle many challenges previously encountered in research  (Zheng et al., 2017; Li et al., 2020a; Paolini et al., 2021; Lu et al., 2022b; Lou et al., 2023; Chen et al., 2022b, 2024), such as adaptability issues when dealing with unseen labels during training. Some studies  (Wei et al., 2023; Wang et al., 2023a; Xie et al., 2023) make significant performance gains in low-resource settings by designing prompt-based frameworks and leveraging models like ChatGPT for in-context learning. Moreover, research efforts such as InstructUIE  (Wang et al., 2023b), PIVOINE  (Lu et al., 2023), and YAYI-UIE  (Xiao et al., 2023), which employ instruction-tuning of open-source LLMs, also achieve notable successes on IE. Additional research explore areas such as prompt learning  (Zhang et al., 2023a), guidelines  (Sainz et al., 2023) and synthetic dataset  (Amalvy et al., 2023). Despite these advancements, current models fine-tuned with an instruction approach face a major challenge: the coarse schema handling strategies in constructing instructions can lead to a mismatch in the number of schema queries between the training and evaluation phases, thus weakening the model’s generalizability.
@@ B.1 Data Collection and Clean
To comprehensively cover various domains and meet the practical demands of IE, we collect IE datasets from multiple sources. IEPile dataset mainly involves bilingual data (Chinese and English) and focuses on three main categories of IE tasks: Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE). The English data mainly comes from the benchmark dataset IEINSTRUCTIONS  (Wang et al., 2023b), while the Chinese data is similar to the Chinese datasets mentioned in the YAYI-UIE  (Xiao et al., 2023). It should be noted that our Chinese dataset collection is conducted concurrently with the aforementioned research.
@@
During the data cleaning process, we address each dataset individually. Firstly, we calculate the text overlap within each dataset’s training, validation, and test sets. If a text is discovered to have multiple occurrences within the same file accompanied by inconsistent annotations, we exclude all corresponding instances from the dataset. Secondly, we compare the text overlap between training, validation, and test sets. If texts from the test set appear previously in the training or validation sets, we would exclude these instances from the training and validation sets. Moreover, we formulate three heuristic rules to eliminate low-quality and meaningless data:
@@ B.2 Schema-Based Instruction Generation
As illustrated in Figure 1, assume that dataset $\mathcal{D}$ possesses a predefined label set $L$. For a given text $S$, the schemas present in its annotation constitute the positive schema set $Pos\_L$, while others form the negative schema set $Neg\_L$. Inspired by the theory of contrastive learning, we construct a hard negative schema dictionary $\mathcal{K}$, where each key represents a unique schema and the associated value is a collection of schemas that are semantically similar to the key schema. Consequently, the set of hard negative schema, $Hard\_L$, is defined as $\mathcal{K}[Pos\_L]$. However, if $Neg\_L$ is composed solely of $Hard\_L$, it would lack a sufficient number of negative instances for the model to learn effectively. Therefore, we define another set of negative schemas, $Other\_L=L-Hard\_L-Pos\_L$. Ultimately, the $Neg\_L$ is composed of $Hard\_L$ and a small number of $Other\_L$ (roughly $split\_num$). The rationale behind the development of these hard negatives is two-fold: firstly, to induce a more frequent co-occur of semantically similar schemas within the instructions, and secondly, to reduce the volume of training instances without sacrificing the model’s performance. Consider a scenario involving 48 schemas paired with a $split\_num$ of 4. Conventionally, this setup would necessitate the generation of 12 distinct instructions. Yet, with the inclusion of hard negatives, this number can be reduced to as few as 3.
@@
0:  Text $S$, Full label set $L$, Positive label set $Pos\_L$, Number of schemas to split $split\_num$
0:  Negative label set $Neg\_L$, Set of instructions $Instructions$ Step 1: Initialize Hard Negative Schema Dictionary $\mathcal{K}$
@@
     $\mathcal{K}[schema]\leftarrow\text{SEMANTIC-CLOSE}(schema,L)$
@@
  $L^{\prime}=Neg\_L\cup Pos\_L$
  Shuffle $L^{\prime}$ to get a randomized sequence
@@
In determining the range for $split\_num$, we combine empirical experimental results with an in-depth analysis of dataset characteristics. For a dataset containing N different labels, the theoretical value of $split\_num$ should fall within the interval [1, N]. Given that we deal with datasets of varying numbers of labels, we aim to identify a $split\_num$ value that could be shared by as many datasets as possible, thereby ensuring that this number is a common divisor for the label numbers of most datasets. For instance, for Named Entity Recognition datasets, we set $split\_num$ to 6; for Relation Extraction and Event Extraction datasets, we establish $split_{n}um$ at 4. We also observe that when $split\_num$ is 1, the ratio of positive to negative samples significantly impacts model performance, and the corresponding number of training samples becomes vast, affecting efficiency adversely. More crucially, we believe that enumerating multiple schemas in instructions aids the model in more effectively learning to distinguish and identify various schemas, thereby enhancing model performance.
Moreover, to enhance model robustness and its clear understanding of the dynamically changing schema sequences in instructions, we set the actual number of schema splits within a dynamic range of [$split\_num$ // 2, $split\_num$ + $split\_num$ // 2]. Specifically, if the number of schemas in the last batch is less than half of $split\_num$, it is merged with the previous batch; it is only treated as a separate batch when its number exceeds half of $split\_num$.
Task Instruction & Output
NER
1{
2    "instruction": "You are an expert in named entity recognition. Please extract entities that match the schema definition from the input. Return an empty list if the entity type does not exist. Please respond in the format of a JSON string.",
3    "schema": ["location", "else", "organization", "person"],
4    "input": "The objective of the Basic Course on War is to provide for combatants of the EPR basic military knowledge for the armed conflict against the police and military apparatus of the bourgeoisie."
5}
6output = {
7    "location": [],
8    "else": [],
9    "organization": ["EPR"],
10    "person": []
11}
RE
1{
2    "instruction": "You are an expert in relationship extraction. Please extract relationship triples that match the schema definition from the input. Return an empty list for relationships that do not exist. Please respond in the format of a JSON string.",
3    "schema": ["place of birth", "country capital", "country of administrative divisions", "company"],
4    "input": "Born on May 1 , 1927 , in Brichevo , Bessarabia in the present-day Republic of Moldova , Mr. Bertini emigrated to Palestine with his family as a child and pursued musical studies there , in Milan , and in Paris , where he worked with Nadia Boulanger and Arthur Honegger."
5}
6output = {
7    "place of birth": [{"head": "Mr. Bertini", "tail": "Paris"}],
8    "country capital": [],
9    "country of administrative divisions": [],
10    "company": []
11}
EE
1{
2    "instruction": "You are an expert in event extraction. Please extract events from the input that conform to the schema definition. Return an empty list for events that do not exist, and return NAN for arguments that do not exist. If an argument has multiple values, please return a list. Respond in the format of a JSON string.",
3    "schema": [{"event_type": "pardon", "trigger": true, "arguments": ["defendant"]}, {"event_type": "extradite", "trigger": true, "arguments": ["person", "agent", "destination", "origin"]}, {"event_type": "sue", "trigger": true, "arguments": ["place", "plaintiff"]}, {"event_type": "start position", "trigger": true, "arguments": ["person", "entity", "place"]}],
4    "input": "Ethical and legal issues in hiring Marinello"
5}
6output = {
7    "pardon": [],
8    "extradite": [],
9    "sue": [],
10    "start position": [{"trigger": "hiring", "arguments": {"person": "Marinello", "entity": "NAN", "place": "NAN"}}]
11}
@@
The instruction format of IEPile adopts a structure akin to JSON strings, essentially constituting a dictionary-type string. This structure is comprised of three main components: (1) “instruction”, which is the task description outlining the objective of the instruction’s execution; (2) “schema”, a list of labels that need to be extracted, specifying the key fields of the information to be extracted; (3) “input”, the source text from which information is to be extracted. Examples of instructions corresponding to various tasks can be found in Table B.2.

### v3

@@ Abstract
Large Language Models (LLMs) demonstrate remarkable potential across various domains; however, they exhibit a significant performance gap in Information Extraction (IE). Note that high-quality instruction data is the vital key for enhancing the specific capabilities of LLMs, while current IE datasets tend to be small in scale, fragmented, and lack standardized schema. To this end, we introduce IEPile, a comprehensive bilingual (English and Chinese) IE instruction corpus, which contains approximately 0.32B tokens. We construct IEPile by collecting and cleaning 33 existing IE datasets, and introduce schema-based instruction generation to unearth a large-scale corpus. Experimentally, IEPile enhance the performance of LLMs for IE, with notable improvements in zero-shot generalization. We open-source the resource and pre-trained models, hoping to provide valuable support to the NLP community.
@@
Honghao Guithanks:     Equal Contribution., Lin Yuan11footnotemark: 1, Hongbin Ye, Ningyu Zhangthanks:     Corresponding Author. Mengshu Sun, Lei Liang, Huajun Chen22footnotemark: 2 Zhejiang University Zhejiang University - Ant Group Joint Laboratory of Knowledge Graph Ant Group {guihonghao,zhangningyu,huajunsir}@zju.edu.cn https://github.com/zjunlp/IEPile
@@ 1 Introduction
By fine-tuning a selection of the latest prominent models (Yang et al., 2023; Touvron et al., 2023b; Bai et al., 2023) on the IEPile dataset, we show that LLMs with IEPile can yield better zero-shot performance than baselines. This achievement not only verifies the effectiveness of the IEPile dataset but also provides a framework for creating IE datasets in other domains.
@@ 2.1 Data Collection and Cleaning
To broadly cover various domains and meet the practical demands, we collect datasets necessary for IE from multiple data sources. Our corpus mainly involves bilingual data (Chinese and English) and focuses on three principal categories of IE tasks: Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE). In total, we gather 26 English datasets and 7 Chinese datasets. We also employ standardization procedures to maintain data quality and format uniformity, involving format unification, instance deduplication, and the exclusion of low-quality data.
@@ 2.2 Schema-Based Instruction Generation
We concentrate on instruction-based information extraction (IE), a methodology that incorporates three crucial elements to compose an instruction: 1) Task Description, a template utilized to distinguish between different IE tasks; 2) Input Text, the source text to be extracted; and 3) Schema sequence, which defines the information that the model is supposed to extract, including entity types, relations, events, etc. Among these, the schema sequence is critical as it reflects the specific extraction requirements and is dynamically variable. Therefore, the construction of the schema sequence within an instruction holds critical significance.
@@
Firstly, we define schemas that actually exist within the input text as positive schemas and those that do not appear as negative schemas. As illustrated in Figure 1, the “location contains” present in the annotation is a positive schema, while all other schemas from the predefined label set $L$ are negative schemas. Traditional IE frameworks, which are treated as sequence labeling tasks, take text as input and produce a label for each token as output, without involving the concept of positive or negative schemas within the model’s input. However, in the era of generative IE, represented by models like UIE  (Lu et al., 2022a), introduce the concept of integrating a schema sequence (refers to as Structural Schema Instructor, or SSI) in the model’s input to guide its output, restricting the range of output to the SSI. The method necessitates including the entire predefined label set of a dataset as the SSI to guide the model’s output during inference. As a result, if the SSI during the training contains only positive schemas, the model will tend to generate corresponding answers for every label within the SSI during inference. Therefore, to make the model explicitly reject generating outputs for negative schemas, it is necessary to incorporate negative schemas into the SSI.
@@
As illustrated in Figure 1, assume that dataset $\mathcal{D}$ possesses a predefined label set $L$. For a given text $S$, the schemas present in its annotation constitute the positive schema set $Pos\_L$, while others form the negative schema set $Neg\_L$. In our analysis, we discover that the primary cause of model mistakes stems from the semantic ambiguity of the schema. In traditional approaches, the $Neg\_L$ is simply defined as $L-Pos\_L$. However, they overlook a critical aspect: it is important to pay special attention to negative schemas that are semantically similar to positive schemas. Inspired by the theory of contrastive learning, we construct a hard negative schema dictionary $\mathcal{K}$, where each key represents a unique schema and the associated value is a collection of schemas that are semantically similar to the key schema. Based on this, we define the hard negative schema set as $Hard\_L=\mathcal{K}[Pos\_L]$, and the other negative schema set as $Other\_L=L-Pos\_L-Hard\_L$. The final $Neg\_L$ is constituted by $Hard\_L$ and a small subset of $Other\_L$. Through this strategy, we not only present semantically similar schemas more frequently within the instruction but also reduce the number of training instances without sacrificing model performance.
@@
Figure 2: Distribution of different tasks, domains, and source datasets within the IEPile.
@@ 2.3 Data Statistics
Based on the aforementioned methods, we obtain the IEPile dataset, which includes roughly 2 million instruction entries and approximately 0.32B tokens (utilizing the Baichuan2 tokenizer). Figure 2 displays the distribution of domains and source datasets within the IEPile, including 33 datasets spanning multiple domains such as general, news, finance, and biomedical. Additionally, Table C.6 provides examples of instructions and outputs for 3 different tasks within the IEPile.
@@
3 Experiments
Based on IEPile, we fine-tune several latest prominent models, then compare their zero-shot generalization capabilities against a range of baseline models. Results of the full supervision evaluation and training details are described in Appendix C.
3.1 Experimental Settings
Evaluation Metrics: We employ span-based Micro-F1 as the metric for measuring model performance. Baselines: We select a range of strong models for comparative analysis, which include UIE, Baichuan2-13B-Chat, GPT-4, InstructUIE, YAYI-UIE, and so on. Zero-shot Benchmark: We collect 13 datasets that are not present in the training set. OneKE: Additionally, we perform full-parameter fine-tuning of the alpaca2-chinese-13b model utilizing IEPile and other proprietary information extraction datasets.
3.2 Main Results
In Tables 1 and  2, we report the zero-shot performance across three tasks and two languages. Overall, after training with the IEPile, the models achieve better results in the majority of tasks. We believe the success is due to the hard negative schema construction and batched instruction generation strategy, which can mitigate the train-eval mismatch and semantic ambiguity for the diverse schema. We also observe that IEPile-models are slightly behind GPT-4 in English NER. We hypothesize that the marginal gap may be attributed to GPT-4’s exposure to a vast corpus of similar data during its training. Moreover, it is essential to note that InstructUIE focuses on English data while IEPile incorporates both English and Chinese data. This disparity in data may influence the capability of the model in English, potentially reducing the performance. Additionally, OneKE achieves the best results in nearly all zero-shot evaluation tasks. We attribute this success to the enhancements brought by full parameter fine-tuning.
3.3 Analysis
Inconsistency in the Number of Schema Queries Hurt Generalization.
@@ 4 Conclusion and Future Work
From the data perspective, our study primarily focuses on schema-based IE, which limits our ability to generalize to human instructions that do not follow our specific format requirements. Additionally, we do not explore the field of Open Information Extraction (Open IE); however, if we remove schema constraints, our dataset would be suitable for Open IE scenarios. Besides, IEPile is confined to data in English and Chinese, and in the future, we hope to include data in more languages. From the model’s perspective, our research evaluates limited models, along with a few baselines due to the computation resources. Theoretically, IEPile can be applied to any other LLMs such as ChatGLM Du et al. (2022) and Gemma  [^3] https://huggingface.co/google/gemma-7b..
@@
Acknowledgements
We would like to express our sincere gratitude to the anonymous reviewers for their thoughtful and constructive feedback. This work was supported by the National Natural Science Foundation of China (No. 62206246), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China (No. LGG22F030011), Yongjiang Talent Introduction Programme (2021A-156-G), CCF-Baidu Open Fund, and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
@@ A.1 Information Extraction Datasets
Large-scale pre-trained corpora are crucial for the effectiveness of LLMs, providing a wealth of knowledge and a foundation for language comprehension. At the same time, the annotated data for information extraction (IE) also holds its importance. Although the field of IE has accumulated a considerable amount of annotated data (Walker et al., 2006; Riedel et al., 2010; Sang and Meulder, 2003; Luan et al., 2018; Gui et al., 2023), these datasets are often limited in size, scattered in distribution, and lack standardization in schema. Faced with these limitations, there is an urgent need for generating instruction data through unified and automated methods to bridge the gap presented by the current absence of centralized, large-scale IE instruction datasets. In this paper, we concentrate on instruction-based IE scenarios. We develop a comprehensive, schema-rich instruction dataset for IE by collecting and cleaning existing IE datasets, called IEPile. IEPile is designed to enhance the adaptability and processing capabilities of LLMs for different IE tasks, simultaneously strengthening their generalization skills to extract from new domains and schemas.
@@ A.2 Information Extraction Models
Recently, LLMs (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023a, b) demonstrate their exceptional versatility and generalization capabilities across a variety of downstream tasks  (Vilar et al., 2023; Hegselmann et al., 2023). Particularly in the domain of IE, these models have the potential to tackle many challenges previously encountered in research  (Zheng et al., 2017; Li et al., 2020a; Paolini et al., 2021; Lu et al., 2022b; Lou et al., 2023; Chen et al., 2022b, 2024), such as adaptability issues when dealing with unseen labels. Some studies  (Wei et al., 2023; Wang et al., 2023a; Xie et al., 2023) make significant performance gains in low-resource settings by designing prompt-based frameworks and leveraging models like ChatGPT for in-context learning. Moreover, research efforts such as InstructUIE  (Wang et al., 2023b), PIVOINE  (Lu et al., 2023), and YAYI-UIE  (Xiao et al., 2023), which employ instruction-tuning of open-source LLMs, also achieve notable successes on IE. Additional research explore areas such as prompt learning  (Zhang et al., 2023a), guidelines  (Sainz et al., 2023) and synthetic dataset  (Amalvy et al., 2023). Despite these advancements, current models fine-tuned with instruction data face a major challenge: the coarse schema handling strategies in constructing instructions could potentially impair the models’ capacity for generalization.
@@ B.1 Data Collection and Clean
To comprehensively cover various domains and meet the practical demands of information extraction (IE), we collect IE datasets from multiple sources. IEPile dataset mainly involves bilingual data (Chinese and English) and three IE tasks: Named Entity Recognition (NER), Relation Extraction (RE), and Event Extraction (EE). The English part mainly comes from the benchmark dataset IEINSTRUCTIONS  (Wang et al., 2023b), while the Chinese data is similar to the Chinese datasets mentioned in the YAYI-UIE  (Xiao et al., 2023). It should be noted that our Chinese dataset collection is conducted concurrently with the aforementioned research.
@@
During the data cleaning process, we address each dataset individually. Firstly, we calculate the text overlap within each dataset’s training, validation, and test sets. If a text is discovered to have multiple occurrences within the same file accompanied by inconsistent annotations, we exclude all corresponding instances from the dataset. Secondly, we compare the text overlap between training, validation, and test sets. If texts from the test set appear previously in the training or validation sets, we would exclude these instances from the training and validation sets. Furthermore, we formulate three heuristic rules to eliminate low-quality and meaningless data:
@@ B.2 Schema-Based Instruction Generation
As illustrated in Figure 1, assume that dataset $\mathcal{D}$ possesses a predefined label set $L$. For a given text $S$, the schemas present in its annotation constitute the positive schema set $Pos\_L$, while others form the negative schema set $Neg\_L$. Inspired by the theory of contrastive learning, we construct a hard negative schema dictionary $\mathcal{K}$, where each key represents a unique schema and the associated value is a collection of schemas that are semantically similar to the key schema. Consequently, the set of hard negative schema, $Hard\_L$, is defined as $\mathcal{K}[Pos\_L]$. However, if $Neg\_L$ is composed solely of $Hard\_L$, it would lack a sufficient number of negative instances for the model to learn effectively. Therefore, we define another set of negative schemas, $Other\_L=L-Hard\_L-Pos\_L$. Ultimately, the $Neg\_L$ is composed of $Hard\_L$ and a small number of $Other\_L$ (roughly $split\_num$). The rationale behind the development of these hard negatives is two-fold: firstly, to induce a more frequent co-occur of semantically similar schemas within the instructions, and secondly, to reduce the volume of training instances without sacrificing the model’s performance. In the context of a dataset comprising 48 schemas with a given $split\_num$ of 4, traditional mode would dictate the creation of 12 unique instructions per data point. However, through the integration of hard negatives, this requisite can be substantially minimized to a mere 3 instructions.
@@
0:  Text $S$, Predefined label set $L$, Positive schema set $Pos\_L$, Number of schemas to split $split\_num$
0:  Set of $Instructions$ Step 1: Initialize Hard Negative Schema Dictionary $\mathcal{K}$
@@
     $\mathcal{K}[schema]\leftarrow\text{SEMANTIC-SIMILAR}(schema,L)$
@@
  $L^{\prime}\leftarrow Neg\_L\cup Pos\_L$
  Shuffle $L^{\prime}$ to obtain a randomized sequence
@@
In the determination of the optimal range for $split\_num$, our methodology integrates empirical results with an in-depth analysis of dataset characteristics. For a dataset containing N different labels, the theoretical value of $split\_num$ should fall within the interval [1, N]. Addressing datasets with heterogeneous label counts, our objective is to identify a $split\_num$ value that offers broad applicability across numerous datasets, thus ensuring this value serves as a common divisor for the majority of dataset label counts. For instance, for Named Entity Recognition datasets, we set $split\_num$ to 6; for Relation Extraction and Event Extraction datasets, we establish $split\_num$ at 4. We also observe that when $split\_num$ is 1, the ratio of positive to negative samples significantly impacts model performance, and the corresponding number of training samples becomes vast, affecting efficiency adversely. More crucially, we believe that enumerating multiple schemas in instructions aids the model in more effectively learning to distinguish and identify various schemas, thereby enhancing model performance.
Furthermore, to enhance model robustness and its clear understanding of the dynamically changing schema sequences in instructions, we set the actual number of schema splits within a dynamic range of [$split\_num$ // 2, $split\_num$ + $split\_num$ // 2]. Specifically, if the number of schemas in the last batch is less than half of $split\_num$, it is merged with the previous batch; otherwise, it stands as an independent batch.
@@
The instruction format of IEPile adopts a structure akin to JSON strings, essentially constituting a dictionary-type string. This structure is comprised of three main components: (1) “instruction”, which is the task description outlining the objective of the instruction’s execution; (2) “schema”, a list of labels that need to be extracted; (3) “input”, the source text from which information is to be extracted. Examples of instructions corresponding to various tasks can be found in Table C.6.